OpenStudy (el_tucan):

find derivative for y = (x^5)(e^(6x^2))... i find this to be awkward

OpenStudy (anonymous):

\[\large y=x^5e^{6x^2}\] By the product rule, you have \[\large y'=\color{red}{(x^5)'e^{6x^2}}+\color{blue}{x^5\left(e^{6x^2}\right)'}\] For the first term, you'd just use power rule. For the second, you'd use chain rule: \[\large(x^5)'e^{6x^2}=\color{red}{5x^4e^{6x^2}}\\ \large x^5\left(e^{6x^2}\right)'=x^5\cdot6(2x)\left(e^{6x^2}\right)=\color{blue}{12x^6e^{6x^2}}\]

OpenStudy (el_tucan):

it's a multiple choice question and that's not one of them...

OpenStudy (anonymous):

The red and blue terms have a common factor of \(\large x^4e^{6x^2}\). Factor that out first.

OpenStudy (anonymous):

You'll find that it's the same as the second answer.

OpenStudy (el_tucan):

oh

OpenStudy (el_tucan):

snoogins, tank you

OpenStudy (anonymous):

You're welcome!